# DEER: A Comprehensive and Reliable Benchmark for Deep-Research Expert Reports

### v1

@@ Appendix A Comparison with Deep Research Benchmarks
In this section, we compare existing deep research benchmarks with our methodology. Table 7 extends and modifies the comparison table proposed in ResearchRubrics ResearchRubrics, and adds new evaluation axes to summarize how DREX differs from prior work. The ResearchRubrics table uses five axes (whether rubrics are human-written, whether experts curate tasks, whether tasks are open-ended, whether non-technical domains are included, and whether an LLM-as-judge is used). On top of this, we add three axes that we consider essential from the perspective of deep research report evaluation: (1) Claim fact-check, (2) Interpretability, and (3) Systematic criteria. Systematic criteria indicate whether the evaluation criteria form a shared hierarchical ontology across tasks, grounded in external expert norms and guidelines.
@@
The first block of the table consists of benchmarks for evaluating search and browsing agents. AcademicBrowse AcademicBrowse, BrowseComp browsecomp, and Mind2Web2 mind2web2 evaluate, respectively, the ability to browse an academic corpus or the open web to produce short answers to complex queries, the ability to perform agentic search across diverse websites, and the consistency between generated answers and cited sources. These benchmarks focus on “how well the system finds information,” i.e., search and browsing strategies, and thus are one step removed from the deep research report quality evaluation that we study.
@@
The second block covers benchmarks that focus on specific component abilities or conceptualizations of deep research. ResearchBench ResearchBench and ResearcherBench ResearcherBench evaluate, respectively, the ability to generate inspiration and hypotheses from papers, and to provide long-form advice on frontier AI research questions based on expert rubrics. Characterizing Deep Research characterizing and LiveDRBench evaluate the recovery of correct claims in search tasks that require many information units and non-trivial reasoning. SPOT SPOT measures how well a system can detect critical errors in published papers, and ExpertLongBench expertlongbench evaluates the ability to generate long-form expert text without external search. These benchmarks finely probe partial abilities such as hypothesis generation, claim discovery, error detection, and expert writing, but do not directly evaluate the quality of full reports produced by web- or literature-based deep research agents.
@@
The third block targets benchmarks that evaluate the quality of deep research reports themselves. ReportBench ReportBench and DeepScholar-Bench DeepScholar-Bench measure report quality in academic survey and related-work settings using citation coverage, overlap with gold surveys, and verifiability of citations. DeepResearchBench deepbench and LiveResearchBench LiveResearchBench evaluate reports for multi-domain web-based deep research tasks using LLM-generated evaluation criteria (RACE), citation-based fact-checking (FACT), and dimensions such as coverage, depth, presentation, and citation accuracy. DeepResearchArena DeepResearchArena automatically constructs large numbers of research tasks from seminar transcripts, and measures factual alignment and overall quality via coverage/conflict/omission (KSR/KCR/KOR) against evidence keypoints extracted from cited URLs and LLM-generated task-specific rubrics (ACE). DeepResearchGym deepgym provides an offline search environment over a public web corpus and an LLM-as-judge protocol, offering a sandbox for evaluating deep research agents’ reports with respect to alignment with information needs, search faithfulness, and overall quality. ResearchRubrics ResearchRubrics provides 101 deep research prompts and over 2,500 human-written rubrics, and is designed so that an LLM-as-judge assigns per-item scores by following these detailed human criteria.
@@
Among these, we analyze DeepResearchBench, LiveResearchBench, and ResearchRubrics in more detail, as they are closest to DREX in terms of evaluation setting. DeepResearchBench and LiveResearchBench are most similar to our web-based deep research report evaluation setting, but both rely heavily on LLMs to define evaluation criteria and assign scores, and their fact-checking is applied only to a limited subset of claims with explicit citations. ResearchRubrics has the advantage that domain experts write 20–40 rubric items for each deep research prompt, enabling fine-grained capture of what constitutes a “good” answer, but it does not perform source-level claim fact-checking, and the rubric structure differs across prompts, making it challenging to interpret system behavior consistently along shared dimensions and rubrics. In contrast, DREX extracts, as exhaustively as possible, all explicit and implicit claims in a report that can be grounded in evidence documents and performs claim-level fact-checking over this set, and evaluates reports using a shared ontology of 7 high-level areas, 25 criteria, and 130 rubric items derived from expert report evaluation norms and guidelines, augmented with task-specific expert guidance. This allows DREX to avoid designing a new rubric from scratch for each task while still applying a dense set of rubric items to each report, and to interpret system performance along standard dimensions and rubric items across tasks.
Table 8: Comparison of Information Verification Pipelines.
@@
Benchmark Explicit Verif. Implicit Verif. Global Context Key Limitation vs. DREX
@@
ReportBench
$\triangle$ Verifies non-cited claims via external web search voting, ignoring internal document grounding.
@@
DeepScholar-Bench
$\triangle$ Relies on physical distance (sliding window $w$); misses long-range semantic dependencies.
DeepResearch Arena
DeepResearchArena
$\triangle$ Evaluates Source $\to$ Report coverage (summarization) rather than Report $\to$ Source verification; misses uncited hallucinations.
LiveDRBench
characterizing
SPOT
SPOT
Evaluates internal error detection against static ground truth; penalizes valid but unannotated error predictions (false positives).
@@ B.1 Topic Domain Analysis
To configure Deep Research tasks (queries), we first analyzed 5045 in-house user queries collected from Deep Research services to ensure a balanced distribution of tasks across domains. The domain distribution of these queries is presented in Figure 4, from which we finalized 8 topic domains, including the top 7 areas and an "Others" category that groups the rest.
@@ B.2 HLE Subject Mapping
Although actual Deep Research service logs are collected in a research-oriented usage context unlike general QA systems, since the majority of users are not limited to experts in specific academic fields, individual queries are not designed with the premise of generating expert-level academic reports or papers immediately. In addition, they include partial queries that depend on the ongoing session context or practical/pragmatic information needs, so using them as-is for evaluation research tasks may lead to a mismatch with the high-difficulty research tasks we aim for in scope and format.
Accordingly, in this study, while maintaining the 8 topic domain composition derived from actual Deep Research logs, we defined subjects from Humanity’s Last Exam (HLE) hle that semantically correspond to each domain and used HLE items belonging to those subjects as seed tasks for each domain. At this time, only items within the difficulty range currently processable by LLMs among all HLE items were selected through a preliminary performance evaluation, and the selected items were remapped to the previously defined Deep Research topic domains to obtain a higher-difficulty academic task composition while preserving the actual usage distribution. The correspondence between each topic domain of the actual Deep Research service and HLE subjects is summarized in Table 9.
B.3 Conversion from HLE QA to DeepResearch Reports
Since the selected HLE items are initially configured in QA format, they are not suitable for deep report generation tasks as they stand. In this study, experts reviewed the question-answer-rationale of each item to identify the core concepts, theories, and phenomena covered, and then reconstructed them into research-oriented task queries suitable for the Deep Research usage context. At this time, deterministic elements such as specific answers, factual conclusions, or proofs presented in HLE were removed from the prompt, and instead, only high-level writing directions such as analysis scope, comparative perspectives, and issues to be discussed were presented to induce the model to develop the logical structure and conclusion on its own.
In addition, if necessary, the basic structure of the output, such as "Introduction–Related Work–Method–Analysis–Discussion," was briefly specified to make the task more similar to the long-form research report generation expected in actual Deep Research use cases. Through this conversion process, simple answer-type QA items were transformed into a report-generation task set with HLE-level high-quality and high-difficulty problem structures, while following the actual Deep Research usage domain distribution.
@@ C.1 Dimension and Criteria Specification
In this section, we detail the 7 dimensions and 26 detailed criteria of the Deep Research Report Evaluation Taxonomy presented in the main text.
Request Completeness
This dimension assesses how systematically and comprehensively the report addresses the user query with respect to the required elements, analysis scope, and final key message. Request Completeness assesses whether all requirements presented in the query are included without omission and developed with sufficient depth. Scope Boundary assesses whether the covered content, excluded items, assumptions, and limitations are clearly defined and consistently maintained. Value Validity assesses whether the final key message encompasses the entire query, is sufficiently specific and practical, and is logically consistent with the evidence and analysis results presented in the report body.
Evidence Validity
@@
This dimension evaluates how accurate and valid the report’s figures and arguments are in terms of calculation, methodology, and logical development. Numeric Accuracy examines whether calculation processes, used formulas/statistical models, and indicators/units are presented without error, are appropriate for the problem context, and are expressed transparently enough for a third party to reproduce and verify. Logical Support evaluates whether the argument is developed consistently with the topic, necessary background/assumptions/inference steps are specified, and major claims and counter-argument responses are persuasively supported without leaps based on facts, data, and interpretations.
Structure Consistency
@@
This dimension evaluates whether the introduction, body, and conclusion, and the structure of each section of the report, are organized consistently with the topic and scope. Introduction checks whether it concisely and clearly presents the topic, problem, significance, and basic scope. The body checks whether the step-by-step argument is developed without omission or deviation, in accordance with the structure and scope presented in the introduction. Conclusion checks whether it synthesizes the body content to complete the purpose of the introduction without introducing new claims or evidence. Section-Level checks whether each section supports the overall structure through clear organizational principles and appropriate connections.
Narration Style
@@
This dimension assesses whether the report faithfully follows the required format and style and is expressed in a way that the reader can read and understand the content without difficulty. The report form checks whether external requirements, such as document length and section system, conform to professional report practices. Writing Quality checks whether sentences are concise and accurate, and whether terminology and tone are consistent. Paragraph checks whether paragraphs are sufficiently developed around a single point and naturally integrated with structural auxiliary elements. Reader Friendliness assesses whether subheadings and simple explanations/examples are used to guide the reader in following complex content.
Ethics Compliance
@@
This dimension assesses whether the report is written ethically and responsibly regarding sensitive issues, potential harm, and perspective balance. Sensitive Issues checks whether sensitive topics such as politics, race, and gender are addressed with neutral, fair language and a balanced perspective. Safety & Impact checks whether negative impacts, side effects, and potential misuse of proposals/technologies/research results are adequately reviewed, and whether specific method presentations are dangerous. Perspective Balance assesses whether the discussion is balanced by appropriately including related perspectives and opposing views without bias toward a specific position.
@@
This dimension assesses whether the information required to answer the request has been adequately secured and presented in terms of quantity and scope. Source Support evaluates whether verifiable evidence or reliable sources are provided to support claims that require external evidence. Information evaluates whether reliable facts and claims are sufficiently presented. Citations evaluate whether sufficient citations are made at necessary points. References are evaluated to determine whether the number of actually used references reaches an appropriate level.
@@
This dimension assesses the factuality of the external information used in the report and the reliability and diversity of the citations/sources supporting it. Claim Factuality measures the proportion of verifiable claims that are determined to be factual. Citation Validity measures the proportion of citations that actually support the claim among citations attached to each claim. Reference Accuracy measures the proportion of sources that correctly support the argument among the presented/used references. Reference Quality assesses whether the sources used are reproducible and reliable. Reference Diversity assesses whether the evidence maintains a sufficient level of source diversity without becoming overly concentrated on a few documents.

### v2

@@ Appendix A Comparison with Deep Research Benchmarks
In this section, we compare existing deep research benchmarks with our methodology. Table 6 extends and modifies the comparison table proposed in ResearchRubrics Sharma et al. (2025), and adds new evaluation axes to summarize how DEER differs from prior work. The ResearchRubrics table uses five axes (whether rubrics are human-written, whether experts curate tasks, whether tasks are open-ended, whether non-technical domains are included, and whether an LLM-as-judge is used). On top of this, we add three axes that we consider essential from the perspective of deep research report evaluation: (1) Claim fact-check, (2) Interpretability, and (3) Systematic criteria. Systematic criteria indicate whether the evaluation criteria form a shared hierarchical ontology across tasks, grounded in external expert norms and guidelines.
@@
The first block of the table consists of benchmarks for evaluating search and browsing agents. AcademicBrowse Zhou et al. (2025), BrowseComp Wei et al. (2025a), and Mind2Web2 Gou et al. (2025) evaluate, respectively, the ability to browse an academic corpus or the open web to produce short answers to complex queries, the ability to perform agentic search across diverse websites, and the consistency between generated answers and cited sources. These benchmarks focus on “how well the system finds information,” i.e., search and browsing strategies, and thus are one step removed from the deep research report quality evaluation that we study.
@@
The second block covers benchmarks that focus on specific component abilities or conceptualizations of deep research. ResearchBench Liu et al. (2025) and ResearcherBench Xu et al. (2025) evaluate, respectively, the ability to generate inspiration and hypotheses from papers, and to provide long-form advice on frontier AI research questions based on expert rubrics. Characterizing Deep Research Java et al. (2025) and LiveDRBench evaluate the recovery of correct claims in search tasks that require many information units and non-trivial reasoning. SPOT Son et al. (2025) measures how well a system can detect critical errors in published papers, and ExpertLongBench Ruan et al. (2025) evaluates the ability to generate long-form expert text without external search. These benchmarks finely probe partial abilities such as hypothesis generation, claim discovery, error detection, and expert writing, but do not directly evaluate the quality of full reports produced by web- or literature-based deep research agents.
@@
The third block targets benchmarks that evaluate the quality of deep research reports themselves. ReportBench Li et al. (2025a) and DeepScholar-Bench Patel et al. (2025) measure report quality in academic survey and related-work settings using citation coverage, overlap with gold surveys, and verifiability of citations. DeepResearchBench Du et al. (2025) and LiveResearchBench Wang et al. (2025) evaluate reports for multi-domain web-based deep research tasks using LLM-generated evaluation criteria (RACE), citation-based fact-checking (FACT), and dimensions such as coverage, depth, presentation, and citation accuracy. DeepResearchArena Wan et al. (2025) automatically constructs large numbers of research tasks from seminar transcripts, and measures factual alignment and overall quality via coverage/conflict/omission (KSR/KCR/KOR) against evidence keypoints extracted from cited URLs and LLM-generated task-specific rubrics (ACE). DeepResearchGym Coelho et al. (2025) provides an offline search environment over a public web corpus and an LLM-as-judge protocol, offering a sandbox for evaluating deep research agents’ reports with respect to alignment with information needs, search faithfulness, and overall quality. ResearchRubrics Sharma et al. (2025) provides 101 deep research prompts and over 2,500 human-written rubrics, and is designed so that an LLM-as-judge assigns per-item scores by following these detailed human criteria.
@@
Among these, we analyze DeepResearchBench, LiveResearchBench, and ResearchRubrics in more detail, as they are closest to DEER in terms of evaluation setting. DeepResearchBench and LiveResearchBench are most similar to our web-based deep research report evaluation setting, but both rely heavily on LLMs to define evaluation criteria and assign scores, and their fact-checking is applied only to a limited subset of claims with explicit citations. ResearchRubrics has the advantage that domain experts write 20–40 rubric items for each deep research prompt, enabling fine-grained capture of what constitutes a “good” answer, but it does not perform source-level claim fact-checking, and the rubric structure differs across prompts, making it challenging to interpret system behavior consistently along shared dimensions and rubrics. In contrast, DEER extracts, as exhaustively as possible, all explicit and implicit claims in a report that can be grounded in evidence documents and performs claim-level fact-checking over this set, and evaluates reports using a shared ontology of 7 high-level areas, 25 criteria, and 101 rubric items derived from expert report evaluation norms and guidelines, augmented with task-specific expert guidance. This allows DEER to avoid designing a new rubric from scratch for each task while still applying a dense set of rubric items to each report, and to interpret system performance along standard dimensions and rubric items across tasks.
@@
Benchmark Explicit Verif. Implicit Verif. Global Context Key Limitation vs. DEER
LiveDRBench
Java et al. (2025)
@@
DeepResearch Arena
Wan et al. (2025)
Evaluates Source $\to$ Report coverage (summarization) rather than Report $\to$ Source verification; misses uncited hallucinations.
LiveResearchBench
Wang et al. (2025)
Checklist- and judge-based evaluation depends on task-specific annotations and LLM judgment; implicit claims and cross-sentence dependencies are not systematically enumerated or verified at the claim level.
SPOT
Son et al. (2025)
$\triangle$ Evaluates internal error detection against static ground truth; penalizes valid but unannotated error predictions (false positives).
@@
Li et al. (2025a)
Verifies non-cited claims via external web search voting, ignoring internal document grounding.
@@
Patel et al. (2025)
Relies on physical distance (sliding window $w$); misses long-range semantic dependencies.
@@ B.1 Topic Domain Analysis
To construct deep research tasks, we analyzed 5,842 in-house user queries collected from our deep research system to estimate real-world domain demand and to guide the benchmark’s target domain distribution. Figure 4 shows the resulting distribution. Based on this analysis, we finalized 11 topic domains: the 10 most frequent domains and an Others category that aggregates all remaining domains.
@@ B.2 HLE Subject Mapping
Although our deep research service logs reflect research-oriented usage rather than general-purpose QA, most users are not domain experts, and their queries are typically not formulated as prompts for expert-level academic reports or papers. Moreover, the logs contain many context-dependent fragments (e.g., follow-up queries within an ongoing session) and pragmatic information needs. Therefore, using these queries directly as evaluation tasks would likely mismatch the high-difficulty, expert-level report-generation setting we target in both scope and format.
Accordingly, we used Humanity’s Last Exam (HLE) Phan et al. (2025) as a source of expert-written, high-difficulty seed items that align with expert-level report generation. To preserve the 11-topic domain composition derived from actual deep research logs, we mapped each domain to one of HLE’s 13 subject domains and sampled HLE items from the corresponding subjects as domain-specific seeds. At this time, we filtered candidates via a preliminary performance evaluation, excluding items that were excessively difficult for LLM-based evaluation and retaining only those within an appropriate difficulty range. As a result, we selected a total of 50 seed items across HLE’s 13 subject domains. The correspondence between deep research topic domains and HLE subject domains is summarized in Table 8.
B.3 Conversion from HLE QA to Deep Research Reports
Because the selected HLE items are presented in a QA format, they are not directly suitable for deep research report-generation tasks in their original form. Accordingly, for each item, a domain expert reviewed the question, answer, and rationale to identify the underlying concepts, theories, and phenomena, and then reformulated it into a research-oriented task query appropriate for the deep research setting. During this reformulation, we removed answer-revealing elements from the prompt, such as specific answers, factual conclusions, and proofs, so that the model must derive the reasoning and conclusions on its own. When necessary, we also included writing guidance that constrains report development—such as the intended scope of analysis, comparative perspectives, and key issues to be addressed—to prevent uncontrolled drift and to enable more fine-grained evaluation of whether required elements are covered. Overall, this conversion reconstructs short-answer QA items into long-form report-generation tasks that require expert-level reasoning and exposition.
Each task query was drafted by one domain expert and cross-reviewed by another expert from the same field. Cross-review was repeated in multiple rounds as needed to check whether the reformulated task (i) is not overly narrow, (ii) requires expert-level domain expertise, (iii) specifies a sufficiently concrete research scope and direction, and (iv) does not contain excessive hints that could steer the model toward the answer or conclusion; the task was revised accordingly. Experts were individuals with a master’s degree in the relevant field or equivalent domain expertise. An example research task is provided in Box 5.
@@ B.3 Conversion from HLE QA to Deep Research Reports
Figure 5: Task Query Example
B.4 Construction of Expert Evaluation Guidance
For each task, we constructed Expert Evaluation Guidance to specify what an expert report for the given prompt must cover. The Guidance includes only the mandatory elements required by the topic (excluding optional content or stylistic preferences) and describes each element in as concrete and verifiable a form as possible so that compliance can be judged. The required elements are derived naturally from the task prompt and its writing requirements, and, when applicable, the Guidance also reflects key concepts implied by the underlying HLE item. In addition, because the prompt’s writing requirements (writing-direction instructions/additional requested constraints) are intended to keep the report from deviating from the intended direction and to enable fine-grained evaluation, they are also included as mandatory elements in the Guidance.
For all 50 tasks, Expert Evaluation Guidance was written in the same expert workflow simultaneously with query reformulation, so that each task’s requirements and evaluation criteria are mutually aligned. As in query reformulation, each Guidance was drafted by one domain expert and cross-reviewed by another expert from the same field. During cross-review, we checked and revised whether (i) the topic’s mandatory content elements were included at a sufficiently detailed level and written in an evaluable way, (ii) optional content or stylistic preferences were not included as mandatory evaluation elements, and (iii) requirements specified in the task prompt’s writing requirements (writing-direction instructions/additional requested constraints) were reflected without omission in the Guidance. An example Expert Evaluation Guidance is provided in Box 6.
Example Expert Evaluation Guidance (MDP / Value Iteration) 1. Formal MDP and Value Iteration Setup
The report precisely defines the components of a finite-state MDP (state space, action space, transition probabilities, reward function, discount factor $\gamma$), states the Bellman optimality equation, and presents the value iteration update rule with consistent notation (e.g., $V_{k+1}=\mathcal{T}V_{k}$), ensuring all subsequent analysis refers unambiguously to this formal framework.
2. Contraction Mapping and Norm Specification
[…]
3. Banach Fixed-Point Theorem Application
The report invokes the Banach fixed-point theorem to establish existence, uniqueness, and geometric convergence of the optimal value function, clearly stating the theorem’s conditions (completeness of space, contraction property) and demonstrating how they are satisfied by the MDP and value iteration operator.
4. Reward Boundedness and Sign Independence
The report analyzes the role of reward boundedness in ensuring geometric convergence, demonstrating that bounded rewards (regardless of sign or symmetry) preserve contraction under the Bellman operator, and proving that unbounded rewards can violate the contraction condition; it explicitly states that the sign or symmetry of the reward interval (e.g., symmetric around zero or non-negative) does not affect convergence as long as boundedness and discounting are maintained.
5. Geometric Convergence Rate Derivation
[…]
6. Edge Case Analysis for Reward Boundedness
The report analyzes representative edge cases for reward boundedness, including symmetric vs. asymmetric intervals, zero rewards, and unbounded or improperly scaled rewards, using theoretical reasoning or controlled simulations to test whether geometric convergence holds or breaks down under each condition.
7. Impact of Reward Sign and Symmetry
The report evaluates whether the sign or symmetry of the reward interval (e.g., symmetric around zero vs. non-negative) affects convergence, proving that geometric convergence depends only on the discount factor and boundedness, not on sign, unless reward shaping alters the effective $\gamma$ or violates boundedness.
8. Numerical Stability and Termination Criteria
The report analyzes practical convergence behavior, including how finite precision arithmetic, reward scaling, and choice of stopping criterion (e.g., $V_{k+1}-V_{k}||_{\infty}<\epsilon$) interact with theoretical guarantees, and demonstrates at least one case where numerical error or poor scaling leads to premature termination or slow apparent convergence.
9. Counterexamples for Non-Convergence
The report constructs at least one verifiable scenario where value iteration fails to converge geometrically—such as when $\gamma=1$, rewards are unbounded, or the contraction condition is violated—and explains how this informs the boundary of the reward range for guaranteed convergence.
10. Discount Factor and Reward Interaction
The report examines how the interplay between $\gamma$ and reward magnitude influences the effective contraction rate, showing that while $\gamma$ controls the rate, reward scaling affects the constant factor in the convergence bound, and that rescaling rewards (e.g., dividing by $R_{\max}$) can normalize behavior across different domains.
11. Illustrative Examples with Reproducible Design
The report includes illustrative examples that feature explicit rules for MDP construction (number of states/actions, transition sparsity, reward assignment), initialization of $V_{0}$, value of $\gamma$, stopping threshold, and random seed control. These examples should be described with enough detail to make the reasoning transparent and verifiable.
12. Limitations and Generalizations
The report discusses limitations of the geometric convergence guarantee, including infinite state spaces, continuous actions, non-stationary environments, or non-linear function approximation, and clarifies whether the reward boundedness condition extends to policy iteration, Q-learning, or other dynamic programming variants, grounding each claim in earlier analysis.
Figure 6: Example of Expert Evaluation Guidance
@@ Appendix C Deep Research Report Evaluation Taxonomy Details
Dimensions Sub-dimensions
Request Fulfillment Completeness, Scope, Helpfulness
Analytical Soundness Quantification, Reasoning
Structural Coherence Introduction, Body, Conclusion, Section
Format & Style Report Format, Writing Quality, Paragraph Quality, Readability
Ethics & Compliance Sensitive Handling, Safety & Impact, Perspective Balance
Information Sufficiency Evidence Coverage, Claim Amount, Citation Amount, Reference Amount
Information Integrity Claim Factuality, Citation Support, Reference Support, Reference Quality, Reference Diversity
Table 9: Deep Research Report Evaluation Taxonomy: 7 Major Dimensions and 25 Sub-dimensions
@@ C.1 Dimension and Criteria Specification
In this section, we detail the 7 dimensions and 25 sub-dimensions of the Deep Research Report Evaluation Taxonomy presented in the main text.
Request Fulfillment
Request Fulfillment evaluates whether the report meets the user’s request at a professional standard. It is assessed along three sub-dimensions: Completeness, Scope, and Helpfulness. Completeness evaluates whether all elements explicitly stated or implicitly required by the query are addressed without omission and with sufficient depth. Scope evaluates whether, in addressing these elements, the report clearly specifies what is included and excluded, as well as its assumptions, constraints, and limitations, and maintains these consistently throughout. Helpfulness evaluates whether the report materially advances the user’s goal by providing information that is sufficiently specific, practical, and actionable for direct use.
Analytical Soundness
@@
Analytical Soundness evaluates how accurate and valid the report’s figures and arguments are in terms of calculation, methodology, and logical development. Quantification examines whether calculation processes, used formulas/statistical models, and indicators/units are presented without error, are appropriate for the problem context, and are expressed transparently enough for a third party to reproduce and verify. Reasoning evaluates whether the argument is developed consistently with the topic, necessary background/assumptions/inference steps are specified, and major claims and counter-argument responses are persuasively supported without leaps based on facts, data, and interpretations.
Structural Coherence
@@
Structural Coherence evaluates whether the introduction, body, and conclusion, and the structure of each section of the report, are organized consistently with the topic and scope. Introduction checks whether it concisely and clearly presents the topic, problem, significance, and basic scope. The Body checks whether the step-by-step argument is developed without omission or deviation, in accordance with the structure and scope presented in the introduction. Conclusion checks whether it synthesizes the body content to complete the purpose of the introduction without introducing new claims or evidence. Section checks whether each section supports the overall structure through clear organizational principles and appropriate connections.
Format & Style
@@
Format & Style assesses whether the report faithfully follows the required format and style and is expressed in a way that the reader can read and understand the content without difficulty. Report Format checks whether external requirements, such as document length and section system, conform to professional report practices. Writing Quality checks whether sentences are concise and accurate, and whether terminology and tone are consistent. Paragraph Quality checks whether paragraphs are sufficiently developed around a single point and naturally integrated with structural auxiliary elements. Readability assesses whether subheadings and simple explanations/examples are used to guide the reader in following complex content.
Ethics & Compliance
@@
Ethics & Compliance assesses whether the report is written ethically and responsibly regarding sensitive issues, potential harm, and perspective balance. Sensitive Handling checks whether sensitive topics such as politics, race, and gender are addressed with neutral, fair language and a balanced perspective. Safety & Impact checks whether negative impacts, side effects, and potential misuse of proposals/technologies/research results are adequately reviewed, and whether specific method presentations are dangerous. Perspective Balance assesses whether the discussion is balanced by appropriately including related perspectives and opposing views without bias toward a specific position.
@@
Information Sufficiency assesses whether the information required to answer the request has been adequately secured and presented in terms of quantity and scope. Evidence Coverage evaluates whether verifiable evidence or reliable sources are provided to support claims that require external evidence. Claim Amount evaluates whether reliable facts and claims are sufficiently presented. Citation Amount evaluates whether sufficient citations are made at necessary points. Reference Amount is evaluated to determine whether the number of actually used references reaches an appropriate level.
@@
Information Integrity assesses the factuality of the external information used in the report and the reliability and diversity of the citations/sources supporting it. Claim Factuality measures the proportion of verifiable claims that are determined to be factual. Citation Support measures the proportion of citations that actually support the claim among citations attached to each claim. Reference Support measures the proportion of sources that correctly support the argument among the presented/used references. Reference Quality assesses whether the sources used are reproducible and reliable. Reference Diversity assesses whether the evidence maintains a sufficient level of source diversity without becoming overly concentrated on a few documents.